OpenStudy (anonymous):

Which equation is not an identity? a. tan(–t) = –tan t b. tan t(csc t + cot t) = sec t + 1 c. sin tcot t = cos t d. cos2t − sin2t = –1

OpenStudy (danjs):

sin^2(x)+cos^2(x) = 1 multiply that by (-1) and check answer d

OpenStudy (danjs):

For C \[\sin(t) \cot(t) = \sin(t)\frac{ \cos(t) }{ \sin(t) } = \cos(t)\] true

OpenStudy (danjs):

the tangent of (-theta) is the same as (-1)tan(theta) by symmetry in the unit circle Draw out both tan(-t) and -tan(t) with an arbitrary angle t, to see how they relate

OpenStudy (anonymous):

I don't understand any of that...

OpenStudy (danjs):